### v1

# Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding

Abstract
Language priors constitute one of the primary causes of hallucinations in Large Vision-Language Models (LVLMs), driving the models to generate linguistically plausible yet visually inconsistent content. The language priors in LVLMs originate from the linguistic knowledge inherited from their pre-trained Large Language Model (LLM) backbone. Consequently, this characteristic is an intrinsic property of the model that remains independent of visual inputs. Inspired by the finding that language priors are consistent across images, we propose Cross-Image Contrastive Decoding (CICD), a simple yet effective training-free method to alleviate language priors in LVLMs. CICD first identifies essential and detrimental priors, and then employs contrastive decoding to eliminate the detrimental ones. This approach simultaneously prevents LVLMs from generating hallucinated content while maintaining textual fluency and coherence. Furthermore, the limited information overlap between images helps prevent visual information loss during contrastive decoding. We validate the effectiveness of CICD on four benchmarks with six LVLMs. Our experiments demonstrate that CICD performs remarkably well in mitigating language priors, especially in the image captioning task, where such priors are most pronounced. Code will be released once accepted.
Index Terms:
Large Vision-Language Model, Multi-Model Hallucination, Language Priors
I Introduction
In recent years, the remarkable success of Large Language Models (LLMs) [1; 2; 3; 4] has provided a viable technological pathway toward Artificial General Intelligence (AGI). Numerous studies [5; 6; 7; 8] have extended these models by incorporating visual inputs, resulting in Large Vision-Language Models (LVLMs) capable of visual comprehension. These models perform various tasks using information from visual inputs, such as image captioning [9; 10; 11; 12], visual question answering [13; 14; 15], and visual grounding [16; 17; 18]. Despite their fluency, LVLMs are prone to hallucinations [19; 20], generating text that contradicts visual evidence.
Hallucinations arise from various factors, most directly from language priors [12; 21; 22; 23], which cause LVLMs to struggle with accurately expressing their understanding of visual information. Since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in the LLMs, resulting in responses that reflect both visual cues and language priors. For example, when describing an office scene, an LVLM may incorrectly generate the word “computer” even if no such object is present in the image.
There are many approaches to alleviating language priors. Some works [24; 25; 12] construct specialized training datasets to reduce models’ reliance on language priors. However, the rapid iteration of model backbones limits the practical applicability of such training-based methods. Recently, a series of studies [21; 22; 23; 26] have used the contrastive decoding algorithm [27] to mitigate language priors, achieving remarkable effectiveness. Specifically, these works construct modified contexts by either perturbing [21; 22] or enhancing [23; 26] the visual features from the original context, and then apply contrastive decoding between the original and modified contexts (i.e., subtracting one logit distribution from the other) to reduce language priors.
These methods are sufficiently simple, training-free, and exhibit excellent generalizability. However, since the modified context is derived from the original context, it retains partial visual information, which can be degraded during contrastive decoding. Moreover, the perturbations introduced may corrupt the model’s features, leading to information distortion.
Figure 1: A case of language prior. Adding the same prefix when captioning another image will leads to the same language priors.
The language priors are derived from LLM backbones and are unrelated to the input image. In other words, the language priors remain consistent across different images, as shown in Fig.1. Inspired by this observation, we propose Cross-Images Contrastive Decoding (CICD), a simple yet highly effective method for mitigating language priors in LVLMs. Concretely, we replace the input image with a different image to create a new context that shares the same language priors. We then subtract this logit distribution from the original one to alleviate the influence of language priors, while preserving the original visual information. Since different images contain limited overlapping visual content, our method effectively prevents visual information loss.
Figure 2: Essential priors and detrimental priors. We random select 500 images from the MSCOCO dataset and visualize LLaVA-1.5’s output by t-SNE algorithm [28]. w/ Image represents regular outputs. w/o Image represents the outputs after removing input image. w/ Neg. Image represents replace the input image by the most dissimilar image in the dataset retrieved by CLIP [29].
Figure 3: A case of JS divergence between the logit distribution of two images.
While language priors contribute to hallucinated content in LVLMs, they also play a role in enhancing output fluency. Therefore, we categorize language priors into two types: (1) essential priors, which are necessary for maintaining text fluency, and (2) detrimental priors, which introduce vision-irrelevant information. As shown in Fig. 2, essential priors demonstrate high consistency across different images, whereas detrimental priors exhibit substantial divergence when generating image-related content. Based on this phenomenon, we distinguish between essential and detrimental priors using cross-image consistency, as illustrated in Fig.3. We observe a distinct divergence in cross-image consistency between essential and detrimental priors. We compute the Jensen-Shannon (JS) divergence between the two logit distributions, and set a threshold to identify essential priors that should be retained.
We validate the effectiveness of our method, CICD, on four benchmarks: CHAIR [10], POPE [14], AMBER [30], and DetailCaps [11]. To evaluate the generalizability of CICD, we conduct experiments on six LVLMs: InstructBLIP [6], LLaVA-1.5 [31], LLaVA-Next [32], Qwen-VL-Chat [7], Qwen2-VL [33], and Qwen2.5-VL [8]. Experimental results demonstrate that CICD effectively mitigates language priors in LVLMs, thereby reducing hallucinations. This improvement is particularly notable in the image captioning task, which typically suffers from stronger language priors due to its longer textual context.
In summary, our main contributions are as follows:
  • We conduct a thorough investigation of language priors and demonstrate that they are intrinsic, model-derived properties independent of visual inputs. Furthermore, we show that consistency across different images can be used to distinguish essential priors from detrimental ones.
  • We propose CICD (Cross-Images Contrastive Decoding), which leverages both the consistency of language priors across images and the divergence of inter-image visual features to mitigate linguistic priors while preserving visual information.
  • Extensive experiments across four benchmarks and six model backbones demonstrate the superiority of our method in alleviating language priors, particularly in the image captioning task.
II Related Work
II-A Large Vision-Language Models
Typically, Large Vision-Language Models (LVLMs) [6; 31; 7; 34] consist of a visual encoder [29; 35] and a Large Language Model (LLM) backbone [2; 3]. The visual encoder extracts visual features, which are then aligned with text tokens through a fusion module to form the input context for the LLM. Through multi-stage training, including modality alignment and instruction tuning [5], LVLMs acquire the ability to perform various multimodal tasks such as image captioning [11], visual question answering [15], and visual grounding [18]. Despite significant advancements, LVLMs still suffer from severe hallucination problems, often generating content that contradicts the visual input.
II-B Hallucination in LVLMs
Originally, hallucinations [36; 37] were defined as undesirable outputs or non-factual information. In the multimodal domain, the study of hallucinations primarily focuses on cross-modal consistency [19; 20], requiring that LVLM-generated content remains faithful to the visual inputs. Multiple factors contribute to hallucinations in LVLMs, including biased visual attention [38; 39], attention misallocation [40; 41; 42], and language priors [21; 23; 22]. Among these, language priors are the most direct cause, making it difficult for LVLMs to accurately express their understanding of visual information.
II-C Language Priors
As LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in LLMs, which are referred to as language priors [24; 25]. Language priors specifically manifest when LVLMs generate content that is linguistically plausible but inconsistent with visual evidence. Some works [12; 43; 44] meticulously design specialized training data to mitigate this issue. However, such training-based approaches exhibit limited generalizability and cannot be easily transferred to broader application scenarios. Recently, a series of training-free approaches have directly reduced language priors using contrastive decoding [27; 45], achieving remarkable performance. These methods construct an alternative logit distribution on top of the original one through techniques such as masking the image [21], perturbing the instruction [22], augmenting the vision input [23], or performing cross-modal conversion [26]. During decoding, the two logit distributions are contrasted to eliminate language priors.
While these methods effectively mitigate linguistic priors, they inevitably incur visual information loss during contrastive decoding due to information overlap, where both distributions share partial visual features. Through a comprehensive analysis of language priors, we identify their vision-independent nature. Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method, which eliminates language priors by leveraging their consistency across different images, while preserving visual information through inter-image variation.
III Preliminary and Motivation
LVLMs retain the core architecture of LLMs while augmenting it with a vision encoder for visual feature extraction and contextual integration. The input image is encoded into visual tokens and inserted into the textual context. After forward propagation, LVLMs project the outputs into a logit distribution to compute conditional probabilities and generate responses following an autoregressive paradigm:
$y_{t}\sim p_{\theta}(y_{t}\mid v,x,y_{<t})=\mathrm{softmax}\left[\mathrm{logit}_{\theta}(y_{t}\mid v,x,y{<t})\right]$ (1)
where $v$ and $x$ denote the visual input and textual input, $y_{t}$ denotes the token generated at the $t$-th step, and $y_{<t}=[y_{1},y_{2},\cdots,y_{t-1}]$ is the sequence of tokens generated before $y_{t}$. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the LVLM with parameters $\theta$, and is normalized into conditional probabilities $p_{\theta}$.
Since LVLMs are fine-tuned from pre-trained LLMs to enhance fluent generation capabilities, they inherently retain prior knowledge that is independent of visual inputs, i.e., language priors. Concretely, tokens ranked higher in the LVLM’s distribution $\mathrm{logit}_{\theta}$ may be semantically plausible but objectively inconsistent with visual evidence. This phenomenon results in hallucinations, where the generated outputs are linguistically fluent yet visually contradictory.
Language priors can be directly suppressed by downweighting the corresponding tokens in the logit distribution. However, since these tokens cannot be explicitly identified, contrastive decoding [27; 21] has emerged as a widely adopted solution to mitigate language priors. This method constructs an auxiliary distribution with enhanced language priors, and then computes a new distribution by contrasting the original and auxiliary distributions:
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ (2)
$-\alpha\,\mathrm{logit}_{C}$
where $\mathrm{logit}_{C}$ is the auxiliary distribution, and $\mathrm{logit}_{\mathrm{CD}}$ is the final distribution used for next-token prediction.
The core principle of contrastive decoding lies in constructing a logit distribution that preserves consistent language priors while excluding consistent visual information. This ensures that visual information is not inadvertently removed along with language priors during contrastive decoding. However, since LVLMs are fine-tuned using multimodal data, removing the visual input distorts the resulting logit distribution, leading to inconsistent language priors (corresponding experiments are presented in Sec.V-G).
We demonstrate that language priors are inherent characteristics of LVLMs and remain consistent across different images, as shown in Fig.1. An intuitive approach is to use other images to construct a logit distribution that shares the same language priors but contains different visual information. Accordingly, Eq.2 is modified as:
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,v^{\prime},x,y_{<t})=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ (3)
$-\alpha\,\mathrm{logit}_{\theta}(y_{t}\mid v^{\prime},x,y_{<t})$
where $v\textquoteright$ ’ is a different image from $v$. $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y{<t})$ encodes both the visual information $V$ from image $v$ and the language priors $T$, which we denote as $V\circ T$. The effectiveness of contrastive decoding can be formulated as:
[…]
where $V\textquoteright$ ’ denotes the visual information of image $v\textquoteright$ ’, and $I(V;V\textquoteright)$ ’ ) denotes the consistency between $V$ and $V\textquoteright$ ’. This approach achieves two objectives simultaneously: (1) precise elimination of language priors, and (2) full preservation of visual information integrity.
IV Method
Leveraging the cross-image consistency of language priors, we propose CICD (Cross-Images Contrastive Decoding), which effectively mitigates language priors while preventing visual information loss. In addition, we retain essential linguistic priors to ensure fluent and natural generation. As illustrated in Fig. 4, CICD comprises three key steps: (1) retrieving a distinct image, (2) identifying detrimental language priors, and (3) removing these priors through contrastive decoding.
Figure 4: The input image and another distinct image are fed into LVLM respectively with identical text-prefix. We calculate the JS discrepancy between the two logit distributions to discriminate essential priors and detrimental priors. For detrimental priors, we perform contrastive decoding to alleviate language priors, while applying regular decoding for essential priors. The final output of this case is shown in the right part.
IV-A Image Retrieval
To prevent visual information loss during contrastive decoding, we select an image that has limited information overlap with the input image. Since visual features are extracted by the LVLM’s vision encoder, an effective strategy is to measure the similarity between two images using the vision encoder to estimate their information overlap. In other words, the image with the lowest similarity to the input image is selected as the reference image. For better generalizability, we adopt the pre-trained CLIP [29] as the retriever. We vectorize all images in the dataset in advance, and during inference, retrieve the one with the minimum cosine similarity to the input image.
Although vector retrieval is time-efficient given the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database. As illustrated in Fig.1, we observe that language priors primarily manifest in content words such as objects, attributes, and relations. These visual elements tend to exhibit minimal information overlap across natural images. As a result, randomly selecting an image can be an effective and lightweight alternative in practical implementations. We further discuss this in Sec.V-G.
To ensure reproducibility, we adopt the retrieval-based approach for deterministic image selection throughout our experiments.
IV-B Identifying Detrimental Priors
Although language priors can lead LVLMs to generate hallucinated content, they are still essential for producing fluent and coherent text. Therefore, we categorize language priors into two types: essential priors and detrimental priors, where only the latter need to be eliminated. We define essential priors as purely linguistic attributes that are entirely independent of visual information. Specifically, for essential priors, the logit distributions produced by LVLMs should remain highly consistent across diverse visual inputs. This intrinsic property enables a clear distinction between essential and detrimental priors.
Figure 5: 500 random images from the MSCOCO dataset captioned by LLaVA-1.5. For each next-token prediction, we statistic the JS divergence (JSD) between two logit distributions produced by different images.
We employ Jensen-Shannon (JS) divergence to measure the consistency between two logit distributions produced by different images:
$\mathcal{D}_{\mathrm{JS}}=\mathrm{JS}\left(p_{\theta}(y_{t}\mid v,x,y_{<t})||p_{\theta}(y_{t}\mid v\textquoteright,x,y_{<t})\right)$ (5)
where $p_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq.1. The statistical results are visualized in Fig.5. We observe that the JSD values can be clearly divided into two distinct ranges: $\log_{10}(\mathcal{D}{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$. Larger JSD values correspond to detrimental priors, while smaller values indicate essential priors. Based on this observation, we set a threshold of $\gamma=-4$ to distinguish between the two types of priors.
IV-C Cross-Images Contrastive Decoding
The model performs regular decoding when essential priors are detected, and applies contrastive decoding to eliminate detrimental priors. For detrimental priors, we hypothesize that smaller distributional differences indicate stronger detrimental effects. Consequently, we define the contrastive coefficient $\alpha$ in Eq. 3 as:
$\alpha=1-\log_{10}(\mathcal{D}_{\mathrm{JS}})$ (6)
Based on the observed density of $\mathcal{D}_{\mathrm{JS}}$ values, we clip $\alpha$ to the interval $[1,3]$ to ensure numerical stability.
The final logit distribution is defined as follows:
$\mathrm{logit}_{CICD}(y_{t}\mid v,v^{\prime},x,y_{<t})=$ (7)
$\begin{cases}\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}\leq\gamma\\ \mathrm{logit}_{CD}(y_{t}\mid v,v^{\prime},x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}>\gamma\end{cases}$
where $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 1, and $\mathrm{logit}_{CD}(y_{t}\mid v,v^{\prime},x,y_{<t})$ is defined in Eq. 3.
V Experiment
V-A Benchmarks
To evaluate the effectiveness of our method in eliminating language priors, we employ four multimodal hallucination benchmarks: two generative benchmarks (CHAIR [10] and DetailCaps [11]), one discriminative benchmark (POPE [14]), and one hybrid benchmark (AMBER [30]).
CHAIR assesses the proportion of hallucinated objects, which are generated by the model but not present in the reference annotations. Following prior works, we randomly select 500 images from the MSCOCO [46] dataset as the test set. We additionally select another 500 images to construct a validation set for hyperparameter tuning.
DetailCaps is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V [47], Gemini-1.5-Pro [48], and GPT-4o [49] for evaluation. It contains 4,870 images from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments.
POPE is a widely adopted benchmark for evaluating object hallucinations by prompting LVLMs to identify whether a specific object is present in the image. It consists of 1,500 images, including 1,000 from MSCOCO and 500 from GQA [50].
AMBER combines generative and discriminative tasks, and is evaluated on a curated set of 1,004 images. In addition to image captioning, it includes 14,216 questions designed to assess hallucinations in object, attribute, and relation recognition.
V-B Evaluated LVLMs
We evaluate our method on six widely-used LVLMs: InstructBLIP [6], LLaVA-1.5 [31], LLaVA-Next [32], Qwen-VL-Chat [7], Qwen2-VL [33], and Qwen2.5-VL [8]. InstructBLIP and LLaVA-1.5 are based on Vicuna-7B [3] as their language backbone, while LLaVA-Next is built upon LLaMA3. The Qwen-VL series models, including Qwen2-VL and Qwen2.5-VL, are developed based on the QwenLM framework. Regarding vision encoders, InstructBLIP, LLaVA-1.5, LLaVA-Next, and Qwen-VL-Chat all use the pretrained CLIP model [29], whereas Qwen2-VL and Qwen2.5-VL employ a fine-tuned ViT [35]. All models are implemented at the 8B parameter scale.
TABLE I: Results on POPE. Acc. and F1 represent the average accuracy and F1-score across three datasets (MSCOCO, A-OKVQA, and GQA), respectively. The bolded results denote the best in each setting and the underlined results denote the second-best.
Setting Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
Acc. $\uparrow$ F1$\uparrow$ Acc. $\uparrow$ F1$\uparrow$ Acc. $\uparrow$ F1$\uparrow$
Random Regular 83.47 82.25 80.49 81.10 82.32 79.35
VCD 85.30 84.59 81.75 82.06 84.74 82.80
ICD 84.81 83.44 80.95 81.59 83.10 80.52
IBD 85.36 84.28 74.82 77.87 83.07 80.23
DeFG 86.68 85.71 84.35 84.50 84.04 81.49
CICD 87.85 87.77 84.00 84.89 87.54 86.37
Popular Regular 79.56 78.99 76.59 78.16 81.11 78.34
VCD 80.81 80.84 77.22 78.60 82.95 81.17
ICD 81.44 80.53 76.82 78.48 81.26 78.87
IBD 81.68 81.13 69.35 74.31 82.20 79.59
DeFG 82.96 82.46 79.25 80.46 83.02 80.57
CICD 84.83 85.18 81.01 82.53 85.94 84.90
Adversarial Regular 76.95 77.13 72.89 75.58 78.23 75.72
VCD 77.27 78.19 74.03 76.39 80.37 78.96
ICD 78.07 77.98 73.19 75.91 79.03 77.05
IBD 78.26 78.52 66.60 72.63 79.68 77.36
DeFG 79.40 79.69 75.19 77.52 80.42 78.24
CICD 78.60 80.30 75.55 78.63 81.94 81.37
V-C Baselines
We compare our method with several training-free baselines that specifically target language priors. VCD [21] introduces noise into the input image to suppress language priors, and then applies contrastive decoding for their elimination. ICD [22] constructs negative logits using adversarial instructions and enhances positive signals through contrastive decoding. IBD [23] strengthens the model’s attention to visual features, thereby increasing the influence of image information during decoding. DeFG [26] converts the generated caption into an image and re-generates the caption from that image, leveraging the preserved visual content to reduce hallucinations.
V-D Implementation Details
We use sampling-based decoding for next-token prediction with default settings. Following [21], we incorporate adaptive plausibility constraints into the decoding process. We use the MSCOCO 2014 validation set as the image retrieval corpus, which contains 40,504 images. For retrieval, we adopt CLIP111openai/clip-vit-large-patch14-336 as the feature extractor and perform similarity-based retrieval using precomputed image embeddings. All experiments are conducted on a single 24GB NVIDIA RTX 3090 GPU.
V-E Main Results
Results on POPE
POPE is a discriminative benchmark that incorporates three distinct datasets: MSCOCO, A-OKVQA, and GQA. Each dataset uses three different negative sampling settings: Random, Popular, and Adversarial. We conducted a comprehensive comparison between our method and other baselines across three distinct LVLMs, as detailed in Tab. I. For streamlined presentation, we report the average performance across all three datasets as the final result.
Our method demonstrates superior overall performance, particularly in the most critical metric, achieving the highest F1 scores across all experimental settings. This suggests that we can utilize cross-image consistency to effectively alleviate language priors and promote LVLMs to convey faithful visual information. Additionally, our method demonstrates consistently superior performance across diverse LVLMs, evidencing remarkable generalization capability. As a training-free approach, it offers immediate deployability in practical applications. Another key finding reveals that our method achieves significantly enhanced performance in the Popular setting, demonstrating an average F1 score improvement of 2.84 over the second-best approach. The Popular setting uses the most frequent objects to construct negative samples, which makes LVLMs more susceptible to language priors. The most substantial performance improvements achieved by our method under this setting further demonstrate its exceptional efficacy in mitigating language priors.
TABLE II: Results on the test set of CHAIR. MaxLen=512 indicates setting the maximum generation length of 512 tokens for the model. The bolded results denote the best in each setting and the underlined results denote the second-best.
Setting Method LLaVA-1.5 InstructBLIP
[…]
MaxLen=512 Regular 54.6 16.4 72.6 106.8 62.6 19.5 66.9 113.6
VCD 59.8 17.8 75.6 104 64.8 18.8 71.9 108.6
ICD 57.0 15.0 74.6 103.3 59.0 17.1 69.2 108.5
IBD 57.6 16.5 74.2 104.3 57.6 15.7 70.8 118.5
DeFG 57.4 16.3 75.5 102.8 59.0 17.7 71.5 108.9
CICD 43.8 11.7 75.0 103.3 49.8 13.7 70.3 98.8
MaxLen=64 Regular 24.4 8.9 56.6 53.7 35.6 13.2 56.4 54.8
VCD 25.0 8.3 59 54.0 32.2 10.3 60.6 55.2
ICD 23.2 8.1 58.4 54.4 29.8 9.8 60.6 55.7
IBD 21.2 6.9 58.8 54.5 27.8 9.2 60.3 56.1
DeFG 22.4 7.2 58.2 54.4 32.4 11 59.4 56.0
CICD 18.0 6.1 59.6 53.3 23.8 7.7 62.2 54.8
TABLE III: Results on AMBER. We set the maximun generation length to 512 for the models in this benchmark. AMBER metirc is calculated by $(100-CHAIR+F1)/2$. The bolded results denote the best in each setting and the underlined results denote the second-best.
Model Method Generative Discriminative
[…]
LLaVA-1.5 Regular 11.6 49.7 47.7 4.4 67.4 83.9 61.8 71.2 79.80
VCD 9.8 51.2 43.8 4.4 68.1 85.1 61.0 71.1 80.65
ICD 8.8 51.2 38.7 4.1 70.3 86.0 64.1 73.4 82.30
IBD 9.8 50.5 42.2 4.4 69.2 86.3 62.1 72.2 81.20
DeFG 9.1 50.7 39.9 4.1 70.2 86.8 63.0 73.0 81.95
CICD 6.6 52.7 34.8 2.2 71.1 89.4 61.9 73.1 83.25
InstructBLIP Regular 12.4 51.9 52.4 5.0 68.3 78.6 71.7 75 81.30
VCD 9.9 54.0 44.6 4.2 70.6 80.6 72.7 76.4 83.25
ICD 9.8 53.9 46.7 5.1 69.9 80.3 70.9 75.3 82.75
IBD 9.0 56.1 45.1 4.6 54.2 86.3 34.2 49.00 70.00
DeFG 9.7 54.1 44.5 5.2 72.1 80.7 75.1 77.8 84.05
CICD 7.1 53.6 35.0 2.3 72.8 85.4 70.1 77.0 84.95
Results on CHAIR
CHAIR is specifically designed to assess object hallucination in the image captioning task, which is one of the most language-prior-susceptible tasks in multimodal research. We evaluate our method in two settings with different maximum generation lengths; the results are shown in Tab. II.
Our method demonstrates superior performance across all experimental conditions, with particularly pronounced improvements observed at longer generation lengths. Longer context lengths amplify the model’s susceptibility to language priors, making these experimental results particularly compelling evidence of our method’s efficacy in mitigating such priors. Compared to baselines, our method’s principal advantage lies in its complete preservation of visual information, which is directly reflected in the superior experimental results. Additionally, we preserve the essential language priors, which are also important in generative tasks.
Results on AMBER
AMBER comprises generative tasks (image captioning) and discriminative tasks (image understanding). The results are shown in Tab. III. We can observe that our method achieves the lowest hallucination frequency in generative tasks while simultaneously attaining state-of-the-art performance on the comprehensive AMBER metric. This superior performance stems from our focused approach to language prior mitigation. This problem is particularly critical for generative tasks, where language priors manifest more severely than in discriminative tasks, explaining our method’s pronounced advantages in image captioning.
TABLE IV: Results on DetailCaps. The bolded results denote the best in each setting and the underlined results denote the second-best.
[…]
LLaVA-1.5 InstructBLIP Qwen-VL-Chat
Regular 52.60 52.99 30.73
VCD 52.91 53.20 31.28
ICD 52.82 53.24 29.70
IBD 52.48 54.14 29.31
DeFG 52.72 53.06 30.21
CICD 55.80 54.20 31.34
Results on DetailCaps
We further validate our method’s effectiveness on the image captioning task, which is more susceptible to language priors. Tab. IV presents the evaluation results on DetailCaps. The CAPTURE metric quantitatively evaluates caption quality through both lexical matching and semantic alignment, measuring the consistency of key content (entities, attributes, and relations) between generated captions and multiple ground-truth references. Our method achieves the best performance across all three evaluated LVLMs, demonstrating its robust capability to mitigate language priors and significantly enhance image captioning performance.
TABLE V: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA). The bolded results denote the best in each setting.
Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
[…]
LLaVA-next Regular 39.0 10.8 62.1 85.47 83.14 79.87
CICD 37.6 10.4 62.7 88.22 86.29 82.76
Qwen2-VL Regular 53.8 10.0 69.4 88.33 85.82 82.43
CICD 35.0 9.2 60.5 90.67 88.07 84.53
Qwen2.5-VL Regular 41.2 11.5 65.7 87.83 85.66 83.12
CICD 37.2 9.3 60.0 89.95 88.00 84.75
V-F Effectiveness on More LVLMs
To comprehensively demonstrate the effectiveness and generalizability of our method, we select three of the most recent open-source LVLMs to conduct experiments. The results are presented in Tab. V. Our method demonstrates significant performance improvements in all experiments, consistently reducing hallucinations across both generative and discriminative tasks through effective mitigation of language priors. The improvements observed across different LVLMs confirm our method’s practical advantages: an elegantly simple framework coupled with state-of-the-art effectiveness.
TABLE VI: This experiment conducted by LLaVA-1.5. Random Image and Retrieved Image represent the contrastive images selected through random sampling and image retrieval, respectively. w/o Image represents the contrastive distribution is constrcuted by removing the input image.
[…]
CHAIR(MaxLen=512)
Regular 54.6 16.4 72.6
w/o Image 54.6 15.5 76.6
Random Image 46.6 13.2 76.7
Retrieved Image 43.8 11.7 75
CHAIR(MaxLen=64)
Regular 24.4 8.9 56.6
w/o Image 25.4 7.6 61.9
Random Image 21.6 6.9 58.8
Retrieved Image 18.0 6.1 59.6
V-G Divergence Between Images
Contrastive decoding is a method that suppresses consistent features between two distributions while amplifying distinctive ones. To prevent the loss of visual information during contrastive decoding, we select an image with sufficient discrepancy to construct a contrasting distribution. Thus, while maintaining language prior consistency, the divergent visual information helps avoid erroneous debiasing caused by feature overlap. We use CLIP to retrieve images based on cosine similarity, and select the image with the lowest similarity as the contrastive image. We conduct experiments to explore the correlation between similarity and performance, as shown in Fig. 6.
We observe that the similarity between images exhibits no strong correlation with performance. Compared to regular decoding methods, images with varying similarity levels all demonstrate significant improvements. As discussed in Sec.IV-A, the influence of language priors is primarily reflected in the LVLMs’ comprehension of objects, attributes, and relationships. This information exhibits limited overlap across real-world images. We hypothesize that random image selection could achieve comparable effectiveness, with experimental results presented in Tab.VI. Experimental results demonstrate that random image selection can also significantly mitigate language priors in LVLMs. In practical applications, the random selection approach offers greater convenience, substantially enhancing the usability of our method.
In Tab. VI, we also attempted to construct language priors by removing visual inputs, but found this approach less effective compared to eliminating linguistic priors through cross-image consistency. We attribute this limitation to LVLMs being trained on large-scale multimodal data, where removing visual inputs disrupts the model’s intrinsic representations and leads to information distortion. Consequently, language priors in LVLMs cannot be directly constructed, which further substantiates the necessity of leveraging cross-image consistency in our method.
Figure 6: The correlation between similarity and performance. We conduct this experiment on the validation set of CHAIR. Sim-CLIP represents normalized image similarity, measured as the cosine distance between the features of two images extracted by CLIP.
Figure 7: The results on the validation set of CHAIR using LLaVA-1.5. The X-axis represents the value of $\gamma$.
V-H Threshold of Detrimental Priors
While language priors contribute to hallucinations in LVLMs, they also play a critical role in maintaining textual coherence and fluency. We employ JS divergence to distinguish between essential and detrimental priors, where essential priors exhibit high cross-image consistency, reflected in lower JS divergence scores. We establish a threshold of $\gamma=-4$ in Eq.7, based on the statistical distribution of JS divergence scores shown in Fig.5.
A comprehensive analysis is shown in Fig. 7. The experimental results align with the statistical analysis, demonstrating optimal performance at $\gamma=-4$. Another notable finding is that preserving essential priors proves meaningful, as it ensures fluent and coherent generation contexts.
VI Conclusion
In this paper, we focus on the language priors issue in LVLMs. Through a comprehensive investigation, we reveal that language priors are intrinsic characteristics of the model and remain consistent across different images. Based on this observation, we propose Cross-Images Contrastive Decoding (CICD)—a training-free method that effectively mitigates language priors while preserving visual information. Extensive experiments across multiple benchmarks and LVLM architectures validate the effectiveness of CICD, particularly in image captioning tasks where language priors are most pronounced.
The simplicity and efficiency of CICD give it substantial practical potential, enabling synergistic integration with other hallucination mitigation methods to further enhance LVLM performance. This constitutes our primary direction for future work.