### v2

# Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding

Abstract
Language priors are a major cause of hallucinations in Large Vision-Language Models (LVLMs), often leading to text that is linguistically plausible but visually inconsistent. Recent work explores contrastive decoding as a training-free solution, but these methods typically construct negative visual contexts from the original image, resulting in visual information loss and distorted distribution. Motivated by the observation that language priors stem from the LLM backbone and remain consistent across images, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective training-free method that uses different images to construct negative visual contexts. We further analyze the cross-image behavior of language priors and introduce a distinction between essential priors (supporting fluency) and detrimental priors (causing hallucinations), enabling selective suppression. By selectively preserving essential priors and suppressing detrimental ones, our method reduces hallucinations while maintaining coherent and fluent language generation. Experiments on four benchmarks and six LVLMs across three model families confirm the effectiveness and generalizability of CICD, especially in image captioning, where language priors are particularly pronounced. Code will be released upon acceptance.
Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding
Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Chong Feng1,3,thanks: Corresponding Autor: chongfeng@bit.edu.cn , 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China, 2Zhongguancun Academy, Beijing, China, 3Southeast Academy of Information Technology, Beijing Institute of Technology, Fujian, China,
1 Introduction
Figure 1: A case of language prior. Adding the same prefix when captioning another image will leads to the same language priors. The model used is LLaVA-1.5.
Recent advances in Large Language Models (LLMs) Brown et al. (2020); Touvron et al. (2023); Chiang et al. (2023); Bai et al. (2023a) have opened promising pathways toward Artificial General Intelligence (AGI). Building on these developments, numerous studies Liu et al. (2023); Dai et al. (2023); Bai et al. (2023b, 2025) have extended LLMs by integrating visual inputs, giving rise to Large Vision-Language Models (LVLMs) capable of understanding visual content. They are capable of addressing a broad spectrum of vision-language tasks, such as image captioning Jing et al. (2024); Rohrbach et al. (2018); Ye et al. (2025); Chen et al. (2025), visual question answering (VQA) Lovenia et al. (2024); Li et al. (2023b); Yu et al. (2024), and visual grounding Rasheed et al. (2024); You et al. (2024); Lu et al. (2024).
Despite impressive capabilities, LVLMs remain susceptible to hallucinations Li et al. (2023c); Wang et al. (2023), often generating text that contradicts the visual input. These hallucinations stem from various sources, with language priors Chen et al. (2025); Leng et al. (2024); Wang et al. (2024c); Zhu et al. (2024b) being the most fundamental underlying cause. Since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge encoded in the LLMs. As a result, their outputs are influenced not only by visual evidence but also by pre-learned language patterns. This often leads to a tendency to rely on memorized linguistic associations when responding to user queries. For instance, when describing an office scene, an LVLM may erroneously mention a "computer" even if no such object is present in the image.
Figure 2: Essential priors (left) and detrimental priors (right). We randomly select 500 images from the MSCOCO dataset and visualize LLaVA-1.5’s output by the t-SNE algorithm van der Maaten and Hinton (2008). w/ Image represents regular outputs. w/o Image represents the outputs after removing the input image. w/ Neg. Image represents replacing the input image by the most dissimilar image in the dataset retrieved by CLIP Radford et al. (2021).
Methods for mitigating language priors in LVLMs can generally be classified into two categories. The first involves constructing specialized training datasets to reduce the model’s dependence on linguistic biases Zhibo et al. (2023); Wu et al. (2022); Chen et al. (2025), though such approaches often incur substantial training costs. In this work, we focus on plug-and-play, training-free methods. Many recent studies adopt contrastive decoding Li et al. (2023a), which constructs negative contexts to expose and suppress language priors. Some approaches directly perturb the input image Leng et al. (2024); Wang et al. (2024c) to alter the visual context, while others enhance the visual features extracted from the original image Zhu et al. (2024b); Zhang et al. (2025). These methods typically compute the difference between the logit distributions obtained from the original and modified (i.e., negative) contexts to reduce the model’s reliance on language priors. Following the paradigm introduced in VCD Leng et al. (2024), these methods routinely modify the original image to generate negative visual contexts. However, this prevailing design choice raises a critical question: Is constructing negative contexts through perturbations of the original image truly the most effective strategy for inducing and mitigating language priors?
We argue that the answer is unequivocally negative. Constructing negative contexts based on the original image can lead to two main issues: (1) Visual information loss, where the negative visual context retains partial visual information, which may be degraded during contrastive decoding. (2) Distorted distribution, where perturbing the model’s input to induce priors may disrupt its internal representations, ultimately resulting in a distorted logit distribution. This analysis raises an important question: How can the above issues be avoided when constructing negative contexts?
In fact, simply taking another image to construct negative context is sufficient. This insight is based on the thought that the language priors guiding step-wise token generation in LVLMs originate from the LLM backbone and are largely independent of the input image. As shown in Fig.1, LVLM produces highly consistent outputs when given completely different images but the same prefix, which strongly supports our hypothesis. Building on this, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective method for mitigating language priors in LVLMs. Instead of modifying the original image, CICD constructs the negative context using a different image that shares the same language priors. Our method avoids the two key issues present in previous approaches. Since the overlap between the original and the new image is limited, our approach avoids introducing critical visual information into the negative context. Moreover, it enables the model to reveal intrinsic priors, maintaining the model’s natural inference process without injecting external bias. The formal analysis of our method is presented in Sec. 3.2.
Figure 3: A case of JS divergence between the logit distributions of two images.
Focusing on language priors, we further explore whether all such priors should be strictly eliminated. We argue that the answer is negative. Language priors can be divided into two categories: (1) essential priors, which are necessary for maintaining textual fluency, and (2) detrimental priors, which introduce information unrelated to the visual input. Clearly, only the detrimental priors should be removed. This raises a critical question: How can we accurately distinguish between the two and selectively eliminate only the detrimental ones? As shown in Fig.2, essential priors exhibit high consistency across different images, whereas detrimental priors show substantial divergence when generating image-related content. Based on this observation, we propose to distinguish essential from detrimental priors using cross-image consistency, as illustrated in Fig.3. Specifically, we compute the Jensen–Shannon (JS) divergence between the logit distributions of different visual contexts, and apply a stable threshold to identify essential priors that should be retained.
We validate the effectiveness of CICD on four widely used benchmarks: CHAIR Rohrbach et al. (2018), POPE Li et al. (2023b), AMBER Wang et al. (2024a), and DetailCaps Ye et al. (2025). To assess the generalizability of our approach, we conduct experiments across six LVLMs from three different model families: InstructBLIP Dai et al. (2023), LLaVA-1.5 Liu et al. (2024c), LLaVA-Next Liu et al. (2024d), Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025). Experimental results show that CICD consistently reduces hallucinations by effectively mitigating language priors in LVLMs, outperforming state-of-the-art methods across all benchmarks and models. The improvement is particularly pronounced in the image captioning task, which tends to suffer more from language priors due to its longer textual outputs.
In summary, our main contributions are as follows:
  • We conduct an in-depth analysis of existing contrastive decoding methods and identify their inherent limitations rooted in the reliance on the original image. We further verify the consistency of language priors across images and propose a stable approach for distinguishing essential priors from detrimental ones.
  • Based on this analysis, we propose CICD (Cross-Images Contrastive Decoding), which leverages retrieved images to induce language priors while selectively preserving essential priors during the mitigation process.
  • Extensive experiments on multiple benchmarks and LVLMs demonstrate the superiority of our method in alleviating language priors, particularly in the image captioning task.
2 Related Work
Due to space limitations, we provide a brief overview here. A more detailed discussion of related work can be found in Sec.A.
2.1 Hallucination in LVLMs
Hallucinations in LVLMs primarily concern cross-modal consistency Li et al. (2023c); Wang et al. (2023), which requires that the generated content remains faithful to the visual input. Among the various causes of hallucination, language priors Zhibo et al. (2023); Wu et al. (2022) are among the most fundamental. These priors originate from the LLM’s pretraining corpus and guide generation based on learned language patterns, which can produce fluent yet visually inconsistent outputs. Several approaches Chen et al. (2025); Liu et al. (2024a); Gunjal et al. (2024) attempt to mitigate language priors by constructing specialized training datasets, but these methods often incur significant additional training costs. Alternatively, recent training-free methods have explored contrastive decoding Li et al. (2023a); Chuang et al. (2024), typically constructing negative visual contexts based on the original image Leng et al. (2024); Wang et al. (2024c); Zhu et al. (2024b); Zhang et al. (2025) in an attempt to elicit and suppress language priors.
In this work, we identify critical limitations of these approaches that stem from their reliance on modifying the original image. Specifically, such modifications may lead to visual information loss, biased language priors, and distribution mismatch. We further conduct an in-depth analysis of the cross-image consistency of language priors and, based on this insight, propose CICD. Additionally, we introduce a method for distinguishing between essential priors and detrimental priors, enabling fine-grained and precise mitigation of language priors in LVLMs.
3 Preliminary
Figure 4: We input the target image and a distinct image into the LVLM separately. During the step-by-step generation of the response, we compute the Jensen-Shannon (JS) divergence between the two logit distributions at each step to distinguish essential priors from detrimental ones. For detrimental priors, we apply contrastive decoding to suppress harmful language biases, while regular decoding is used for essential priors.
3.1 Contrastive Decoding
LVLMs take as input a concatenation of vision tokens from a vision encoder and an encoded text prompt. After forward propagation, the model projects the output into a logit distribution to compute conditional probabilities and generate responses in an autoregressive manner:
$y_{t}\sim p_{\theta}(y_{t}\mid v,x,y_{<t})$ (1)
$=$ $\ \mathrm{softmax}\left[\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})\right]$
Here, $v$ and $x$ represent the visual and textual inputs, respectively. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the LVLM with parameters $\theta$.
Since LVLMs adopt pretrained LLMs as their backbone, they inherently inherit language priors that are independent of the visual input. As a result, tokens ranked higher in the output distribution $\mathrm{logit}_{\theta}$ may appear semantically plausible but can be objectively inconsistent with the visual evidence. This constitutes a fundamental cause of hallucinations. A commonly used approach to mitigate such priors is contrastive decoding Li et al. (2023a); Leng et al. (2024), which modifies the original image to construct a native visual context that elicits language priors. These priors are then suppressed by subtracting the logit distribution derived from the native context from that of the original context.
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})$ (2)
$=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})-\alpha\,\mathrm{logit}_{N}$
Here, $\mathrm{logit}_{N}$ denotes the logit distribution obtained from the native context, and $\mathrm{logit}_{\mathrm{CD}}$ represents the final distribution used for next-token prediction.
3.2 Formal Analysis
The core component of contrastive decoding lies in the construction of the native context. Our preliminary investigation reveals that existing methods, which modify the original image, introduce several limitations (see Sec.1). Further observations and analysis show that language priors are inherent characteristics of LVLMs and remain consistent across different images, as illustrated in Fig.1. This insight leads to a natural alternative, where the native context is constructed using an image distinct from the original. Under this design, Eq. 2 is refined as:
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})$ (3)
$=$ $\ (1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$
$\quad-\alpha\,\mathrm{logit}_{\theta}(y_{t}\mid v^{\prime},x,y_{<t})$
where $v\textquoteright$ ’ is a different image from $v$. We further provide a formal analysis of the effectiveness of this approach.
Most existing methods employ sampling decoding, where a token is randomly selected according to the logit distribution. We define the set of plausible tokens under a given logit distribution as $S=V\cup T$, where $V$ represents tokens related to visual content, and $T$ represents tokens unrelated to the image, i.e., those driven by language priors. The negative context constructed by a different image yields a new logit distribution, whose plausible token set is $S\textquoteright=V\textquoteright\cup T$, given that language priors remain unchanged across images. The contrastive decoding operation can thus be interpreted as subtracting one plausible token set from another:
$S-S^{\prime}=(V\cup T)-(V^{\prime}\cup T)=V\mbox{ if }V^{\prime}\cap V=\emptyset$ (4)
In other words, this subtraction removes all language priors that are unrelated to the visual input. This approach achieves three objectives simultaneously: (1) retention of all essential visual information, (2) precise elimination of language priors, and (3) compatibility with the input distribution expected by LVLMs.
4 Method
Leveraging the cross-image consistency of language priors, we propose CICD (Cross-Images Contrastive Decoding), which effectively mitigates language priors while preventing visual information loss. In addition, we retain essential linguistic priors to ensure fluent and natural generation. As illustrated in Fig. 4, CICD comprises three key steps: (1) retrieving a distinct image, (2) identifying detrimental language priors, and (3) removing these priors through contrastive decoding.
4.1 Image Retrieval
As discussed earlier, it is essential to select an image with limited information that overlaps with the input image. Since visual features are extracted by the vision encoder, a natural strategy is to measure the similarity between two images using the encoder itself to estimate their overlap. In other words, the image with the lowest similarity to the input image is selected as the reference. To ensure better generalizability, we adopt the pretrained CLIP model Radford et al. (2021) as the retriever. All images in the dataset are vectorized in advance, and during inference, the image with the lowest cosine similarity to the input is retrieved as the contrastive image.
Although vector retrieval is relatively efficient given the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database. As illustrated in Fig.1, we observe that language priors primarily manifest in content words such as objects, attributes, and relationships. These visual elements tend to exhibit minimal overlap across natural images. Therefore, randomly selecting an image can serve as an effective and lightweight alternative for practical applications. We provide further discussion in Sec.C.2. Note that, to ensure reproducibility, we adopt the retrieval-based strategy for consistent and deterministic image selection throughout all experiments.
4.2 Identifying Detrimental Priors
Although language priors may cause LVLMs to generate hallucinated content, they remain essential for producing fluent and coherent text. To address this, we categorize language priors into two types: essential priors and detrimental priors, with only the latter requiring elimination. Essential priors are defined as purely linguistic patterns that are entirely independent of visual input. As discussed in Sec. 1, the logit distributions associated with essential priors should remain highly consistent across diverse visual inputs. This intrinsic property provides a reliable basis for distinguishing between essential and detrimental priors.
Figure 5: 500 random images from the MSCOCO dataset captioned by LLaVA-1.5. For each next-token prediction, we statistic the JS divergence (JSD) between two logit distributions produced by different images.
We employ the Jensen–Shannon (JS) divergence to measure the consistency between two logit distributions produced from different images:
$\mathcal{D}_{\mathrm{JS}}=\mathrm{JS}\left(p_{\theta}(y_{t}\mid v,x,y_{<t})||p_{\theta}(y_{t}\mid v\textquoteright,x,y_{<t})\right)$ (5)
where $p_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq.1. The statistical results are visualized in Fig.5. We observe that the JSD values can be clearly separated into two distinct ranges: $\log_{10}(\mathcal{D}{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$. Higher JSD values are associated with detrimental priors, while lower values correspond to essential priors, as shown in Fig.2 and Fig.3. Based on this observation, we set a threshold of $\gamma=-4$ to distinguish between the two types of priors.
Setting Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
Acc. $\uparrow$ F1$\uparrow$ Acc. $\uparrow$ F1$\uparrow$ Acc. $\uparrow$ F1$\uparrow$
Random Regular 83.47 82.25 80.49 81.10 82.32 79.35
VCD 85.30 84.59 81.75 82.06 84.74 82.80
ICD 84.81 83.44 80.95 81.59 83.10 80.52
IBD 85.36 84.28 74.82 77.87 83.07 80.23
DeFG 86.68 85.71 84.35 84.50 84.04 81.49
CICD 87.85 87.77 84.00 84.89 87.54 86.37
Popular Regular 79.56 78.99 76.59 78.16 81.11 78.34
VCD 80.81 80.84 77.22 78.60 82.95 81.17
ICD 81.44 80.53 76.82 78.48 81.26 78.87
IBD 81.68 81.13 69.35 74.31 82.20 79.59
DeFG 82.96 82.46 79.25 80.46 83.02 80.57
CICD 84.83 85.18 81.01 82.53 85.94 84.90
Adversarial Regular 76.95 77.13 72.89 75.58 78.23 75.72
VCD 77.27 78.19 74.03 76.39 80.37 78.96
ICD 78.07 77.98 73.19 75.91 79.03 77.05
IBD 78.26 78.52 66.60 72.63 79.68 77.36
DeFG 79.40 79.69 75.19 77.52 80.42 78.24
CICD 78.60 80.30 75.55 78.63 81.94 81.37
Table 1: Results on POPE. Acc. and F1 represent the average accuracy and F1-score across three datasets (MSCOCO, A-OKVQA, and GQA), respectively. The bolded results denote the best in each setting, and the underlined results denote the second-best.
4.3 Cross-Images Contrastive Decoding
After completing the above preparation, we integrate the proposed method into the contrastive decoding process. Specifically, the model performs regular decoding when essential priors are detected and applies contrastive decoding to suppress detrimental priors. In the case of detrimental priors, smaller distributional differences imply a stronger detrimental effect. Accordingly, we define the contrastive coefficient $\alpha$ in Eq. 3 as: $\alpha=1-\log_{10}(\mathcal{D}_{\mathrm{JS}})$. Based on the observed density of $\mathcal{D}_{\mathrm{JS}}$ values, we clip $\alpha$ to the interval $[1,3]$ to ensure numerical stability.
The final logit distribution is defined as follows:
$\mathrm{logit}_{CICD}(y_{t}\mid v,x,y_{<t})=$ (6)
$\begin{cases}\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}\leq\gamma\\ \mathrm{logit}_{CD}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}>\gamma\end{cases}$
where $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 1, and $\mathrm{logit}_{CD}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 3.
5 Experiment
MaxLen Method LLaVA-1.5 InstructBLIP
[…]
64 Regular 24.4 8.9 56.6 53.7 35.6 13.2 56.4 54.8
VCD 25.0 8.3 59 54.0 32.2 10.3 60.6 55.2
ICD 23.2 8.1 58.4 54.4 29.8 9.8 60.6 55.7
IBD 21.2 6.9 58.8 54.5 27.8 9.2 60.3 56.1
DeFG 22.4 7.2 58.2 54.4 32.4 11 59.4 56.0
CICD 18.0 6.1 59.6 53.3 23.8 7.7 62.2 54.8
512 Regular 54.6 16.4 72.6 106.8 62.6 19.5 66.9 113.6
VCD 59.8 17.8 75.6 104 64.8 18.8 71.9 108.6
ICD 57.0 15.0 74.6 103.3 59.0 17.1 69.2 108.5
IBD 57.6 16.5 74.2 104.3 57.6 15.7 70.8 118.5
DeFG 57.4 16.3 75.5 102.8 59.0 17.7 71.5 108.9
CICD 43.8 11.7 75.0 103.3 49.8 13.7 70.3 98.8
Table 2: Results on the test set of CHAIR. MaxLen denotes the maximum generation length.
5.1 Experiment Settings
Due to space constraints, we present only the key aspects of our experimental setup here. Detailed settings can be found in Sec.B.
Benchmarks
We evaluate the performance of CICD across four widely adopted multimodal hallucination benchmarks to validate its effectiveness on different tasks. These include two generative benchmarks (CHAIR Rohrbach et al. (2018) and DetailCaps Ye et al. (2025)), one discriminative benchmark (POPE Li et al. (2023b)), and one hybrid benchmark (AMBER Wang et al. (2024a)).
Evaluated LVLMs
We evaluate the generalizability of CICD across six LVLMs from three model families: InstructBLIP Dai et al. (2023), the LLaVA series (LLaVA-1.5 Liu et al. (2024c) and LLaVA-Next Liu et al. (2024d)), and the Qwen series (Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025)). Unless stated otherwise, all models are implemented at the 7B scale.
Baselines
We compare CICD with four recent training-free methods that specifically target language priors: VCD Leng et al. (2024), ICD Wang et al. (2024c), IBD Zhu et al. (2024b), and DeFG Zhang et al. (2025). All of these methods construct negative visual contexts based on the original image. In contrast, we identify several limitations of this approach and propose CICD as a more effective alternative.
5.2 Implementation Details
We use sampling decoding for next-token prediction with default settings. Following Leng et al. (2024), we incorporate adaptive plausibility constraints into the decoding process. We use the MSCOCO 2014 validation set as the image retrieval corpus, which contains 40,504 images. All experiments are conducted on an NVIDIA RTX 3090 GPU.
Model Method Generative Discriminative
[…]
LLaVA-1.5 Regular 11.6 49.7 47.7 4.4 67.4 83.9 61.8 71.2 79.80
VCD 9.8 51.2 43.8 4.4 68.1 85.1 61.0 71.1 80.65
ICD 8.8 51.2 38.7 4.1 70.3 86.0 64.1 73.4 82.30
IBD 9.8 50.5 42.2 4.4 69.2 86.3 62.1 72.2 81.20
DeFG 9.1 50.7 39.9 4.1 70.2 86.8 63.0 73.0 81.95
CICD 6.6 52.7 34.8 2.2 71.1 89.4 61.9 73.1 83.25
InstructBLIP Regular 12.4 51.9 52.4 5.0 68.3 78.6 71.7 75 81.30
VCD 9.9 54.0 44.6 4.2 70.6 80.6 72.7 76.4 83.25
ICD 9.8 53.9 46.7 5.1 69.9 80.3 70.9 75.3 82.75
IBD 9.0 56.1 45.1 4.6 54.2 86.3 34.2 49.0 70.00
DeFG 9.7 54.1 44.5 5.2 72.1 80.7 75.1 77.8 84.05
CICD 7.1 53.6 35.0 2.3 72.8 85.4 70.1 77.0 84.95
Table 3: Results on AMBER. The maximum generation length is 512. AMBER metirc is calculated by $(100-CHAIR+F1)/2$.
5.3 Main Results
Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
[…]
Regular 55.7 17.4 52.60 58.6 18 52.99 9.9 7.3 30.73
VCD 55.7 16.8 52.91 59.6 18.9 53.20 7.3 5.9 31.28
ICD 53.9 16.6 52.82 55.7 16.7 53.24 5.6 7 29.70
IBD 54.1 15.8 52.48 56.4 15.5 54.14 4.6 2.9 29.31
DeFG 55.7 16.6 52.72 59.1 17.4 53.06 5.6 3.9 30.21
CICD 45.6 13.1 55.80 45.7 13.2 54.20 5.9 3.5 31.34
Table 4: Results on DetailCaps. The maximum generation length is 512.
Results on POPE
As shown in Tab. 1, our method achieves the highest F1 scores across all experimental settings. This suggests that we can utilize cross-image consistency to effectively alleviate language priors and promote LVLMs to convey faithful visual information. The Popular setting uses frequent objects for negative samples, increasing susceptibility to language priors. The strong performance gains of our method in this setting highlight its effectiveness in mitigating such priors.
Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
[…]
LLaVA-Next Regular 41.8 11.9 60.2 85.47 83.14 79.87
CICD 37.6 10.4 62.7 88.22 86.29 82.76
Qwen2-VL-Instruct Regular 53.8 10.0 69.4 88.33 85.82 82.43
CICD 50.6 9.4 73.6 90.67 88.07 84.53
Qwen2.5-VL-Instruct Regular 41.2 11.5 65.7 87.83 85.66 83.12
CICD 39.4 9.6 68.1 89.95 88.00 84.75
Table 5: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA). The bolded results denote the best in each setting.
Results on CHAIR
The results in Tab. 2 demonstrate our method’s superior performance across all experimental conditions, with particularly pronounced improvements observed at longer generation lengths. Longer context lengths amplify the model’s susceptibility to language priors, making these experimental results particularly compelling evidence of our method’s efficacy in mitigating such priors. Compared to baselines, our method’s principal advantage lies in its complete preservation of visual information and essential language priors, which is directly reflected in the superior performance in image captioning.
Results on AMBER
As illustrated in Tab. 3, our method achieves the lowest hallucination frequency in generative tasks while simultaneously attaining state-of-the-art performance on the comprehensive AMBER metric. Language prior is particularly critical for generative tasks, where language priors manifest more severely than in discriminative tasks, explaining our method’s pronounced advantages in image captioning.
Results on DetailCaps
Tab. 4 presents the evaluation results on DetailCaps. The CAPTURE metric quantitatively evaluates caption quality through both lexical matching and semantic alignment, measuring the consistency of key content (entities, attributes, and relations) between generated captions and multiple ground-truth references. Our method achieves the best performance across all three evaluated LVLMs, demonstrating its robust capability to mitigate language priors and significantly enhance image captioning performance.
5.4 Effectiveness on More LVLMs
To comprehensively demonstrate the effectiveness and generalizability of our method, we select three of the most recent open-source LVLMs to conduct experiments. The results are presented in Tab. 5. Our method demonstrates significant performance improvements in all experiments, consistently reducing hallucinations across both generative and discriminative tasks through effective mitigation of language priors. We also test our method on these LVLMs by DetailCaps and the results are shown in Tab. 7. The improvements observed across different LVLMs confirm our method’s practical advantages: an elegantly simple framework coupled with state-of-the-art effectiveness.
Further experiments will be discussed in Appendix B.
6 Conclusion
In this paper, we address the issue of language priors in LVLMs. Through a comprehensive investigation, we reveal that language priors are intrinsic to the model and remain consistent across different images. Building on this observation, we propose Cross-Image Contrastive Decoding (CICD) — a training-free method that effectively mitigates language priors while preserving visual information. Extensive experiments across multiple benchmarks and LVLM architectures demonstrate the effectiveness of CICD, particularly in image captioning tasks where language priors are most prominent.
The simplicity and efficiency of CICD offer strong practical potential, enabling seamless integration with other hallucination mitigation methods to further enhance LVLM performance. This integration forms a key direction for our future work.
7 Limitation
We observe that language priors remain consistent across different images. Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method to eliminate language priors. Our method has two main limitations compared to regular decoding. First, it requires an additional image that differs sufficiently from the original input. Second, both the original and the additional images must undergo separate forward passes to obtain their respective logit distributions for contrastive decoding.
Addtional Image
While the additional image is necessary, there is no strict requirement on its degree of difference from the original. Hallucinations typically occur in content-rich elements such as objects, attributes, and relationships, which naturally exhibit substantial variability across real-world images. Furthermore, our experiments demonstrate that the effectiveness of hallucination mitigation is not strongly correlated with image similarity. The experimental results demonstrate that randomly selected images perform comparably to retrieved ones, highlighting the practical applicability of our method. This flexibility significantly lowers the barrier for real-world deployment.
Twice Forward Propagation
Although performing twice forward propagations doubles the theoretical computational cost, in practice, we merge the two inputs into a single batch and perform one forward propagation with a doubled batch size to achieve the same effect. Under this implementation, our method incurs only a 10% increase in inference time and a 5% increase in memory usage, demonstrating strong potential for practical deployment.
Appendix A Detailed Related Work
A.1 Large Vision-Language Models
Typically, Large Vision-Language Models (LVLMs) Dai et al. (2023); Liu et al. (2024c); Bai et al. (2023b); Zhu et al. (2024a) consist of a visual encoder Radford et al. (2021); Dosovitskiy et al. (2021) and a Large Language Model (LLM) backbone Touvron et al. (2023); Chiang et al. (2023). The visual encoder extracts visual features, which are then aligned with text tokens through a fusion module to form the input context for the LLM. Through multi-stage training, including modality alignment and instruction tuning Liu et al. (2023), LVLMs acquire the ability to perform various multimodal tasks such as image captioning Ye et al. (2025), visual question answering Yu et al. (2024), and visual grounding Lu et al. (2024). Despite significant advancements, LVLMs still suffer from severe hallucination problems, often generating content that contradicts the visual input.
A.2 Hallucination in LVLMs
Originally, hallucinations Ji et al. (2023); Liu et al. (2024b) were defined as undesirable outputs or non-factual information. In the multimodal domain, the study of hallucinations primarily focuses on cross-modal consistency Li et al. (2023c); Wang et al. (2023), requiring that LVLM-generated content remains faithful to the visual inputs. Multiple factors contribute to hallucinations in LVLMs, including biased visual attention Zhu et al. (2025); Li et al. (2025), attention misallocation Huang et al. (2024); Yin et al. (2025); Kang et al. (2025), and language priors Leng et al. (2024); Zhu et al. (2024b); Wang et al. (2024c). Among these, language priors are the most direct cause, making it difficult for LVLMs to accurately express their understanding of visual information.
A.3 Language Priors
As LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in LLMs, which are referred to as language priors Zhibo et al. (2023); Wu et al. (2022). Language priors specifically manifest when LVLMs generate content that is linguistically plausible but inconsistent with visual evidence. Some works Chen et al. (2025); Liu et al. (2024a); Gunjal et al. (2024) meticulously design specialized training data to mitigate this issue. However, such training-based approaches exhibit limited generalizability and cannot be easily transferred to broader application scenarios. Recently, a series of training-free approaches have directly reduced language priors using contrastive decoding Li et al. (2023a); Chuang et al. (2024), achieving remarkable performance. These methods construct an alternative logit distribution on top of the original one through techniques such as masking the image Leng et al. (2024), perturbing the instruction Wang et al. (2024c), augmenting the vision input Zhu et al. (2024b), or performing cross-modal conversion Zhang et al. (2025). During decoding, the two logit distributions are contrasted to eliminate language priors.
While these methods effectively mitigate linguistic priors, they inevitably incur visual information loss during contrastive decoding due to information overlap, where both distributions share partial visual features. Through a comprehensive analysis of language priors, we identify their vision-independent nature. Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method, which eliminates language priors by leveraging their consistency across different images, while preserving visual information through inter-image variation.
Appendix B Detailed Experiment Settings
B.1 Benchmarks
In order to validate the effectiveness of our method in mitigating language hallucination, we evaluate its performance on four widely-used multimodal hallucination benchmarks: two generative benchmarks (CHAIR Rohrbach et al. (2018) and DetailCaps Ye et al. (2025)), one discriminative benchmark (POPE Li et al. (2023b)), and one hybrid benchmark (AMBER Wang et al. (2024a)).
CHAIR
assesses the proportion of hallucinated objects, which are generated by the model but not present in the reference annotations. Following prior works, we randomly select 500 images from the MSCOCO Lin et al. (2014) dataset as the test set. We additionally select another 500 images to construct a validation set for hyperparameter tuning.
DetailCaps
is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V Yang et al. (2023), Gemini-1.5-Pro Team et al. (2024), and GPT-4o OpenAI et al. (2024) for evaluation. It contains 4,870 images from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments.
POPE
is a widely adopted benchmark for evaluating object hallucinations by prompting LVLMs to identify whether a specific object is present in the image. POPE incorporates three distinct datasets: MSCOCO, A-OKVQA, and GQA. Each dataset uses three different negative sampling settings: Random, Popular, and Adversarial. Finally, it consists of 9,000 questions and 1,500 images.
AMBER
combines generative and discriminative tasks, and is evaluated on a curated set of 1,004 images. In addition to image captioning, it includes 14,216 questions designed to assess hallucinations in object, attribute, and relation recognition.
B.2 Evaluated LVLMs
To demonstrate the generalizability of the proposed CICD as a broadly applicable plug-and-play module, we evaluate its performance across six different LVLMs spanning three model families: InstructBLIP Dai et al. (2023), LLaVA-1.5 Liu et al. (2024c), LLaVA-Next Liu et al. (2024d), Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025). InstructBLIP and LLaVA-1.5 are both built upon Vicuna-7B Chiang et al. (2023) as their language backbone, while LLaVA-Next is based on LLaMA3. The Qwen-VL series, including Qwen2-VL-Instruct and Qwen2.5-VL-Instruct, is developed using the QwenLM framework. As for vision encoders, InstructBLIP, LLaVA-1.5, LLaVA-Next, and Qwen-VL-Chat utilize the pretrained CLIP model Radford et al. (2021), whereas Qwen2-VL-Instruct and Qwen2.5-VL-Instruct adopt a fine-tuned ViT Dosovitskiy et al. (2021). All models are implemented at the 8B parameter scale.
B.3 Baselines
We compare our method with several training-free baselines that specifically target language priors.
VCD Leng et al. (2024) introduces noise into the input image to suppress language priors, and then applies contrastive decoding for their elimination.
ICD Wang et al. (2024c) constructs negative logits using adversarial instructions and enhances positive signals through contrastive decoding.
IBD Zhu et al. (2024b) strengthens the model’s attention to visual features, thereby increasing the influence of image information during decoding.
DeFG Zhang et al. (2025) converts the generated caption into an image and re-generates the caption from that image, leveraging the preserved visual content to reduce hallucinations.
Appendix C More Experiment Results
C.1 Effectiveness on Larger LVLMs
We evaluate the effectiveness of our method on larger models using two benchmarks, POPE and CHAIR, with the results presented in Tab.6. Our method continues to deliver significant performance improvements on a 13B model, indicating its strong generalizability. These results suggest that our approach scales well and remains effective across different model sizes.
Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
[…]
LLaVA-1.5-13B Regular 58.8 17 73.4 82.94 81.46 78.49
CICD 44.4 12.9 76.0 87.92 85.68 80.99
INstructBLIP-13B Regular 65.6 21.0 62.7 81.31 77.20 74.69
CICD 44 11.6 65.7 87.52 83.60 78.96
Table 6: Results on larger LVLMs. The maximum generation length is 512.
Method LLaVA-Next Qwen2-VL-Instruct Qwen2.5-VL-Instruct
[…]
Regular 61.06 38.0 10.4 60.04 51.3 10.6 62.62 42.2 11.7
CICD 62.19 36.2 8.5 61.79 48.4 9.9 65.07 43.8 10.9
Table 7: Results on DetailCaps with more LVLMs.
[…]
CHAIR(MaxLen=512)
Regular 54.6 16.4 72.6
w/o Image 54.6 15.5 76.6
Random Image 46.6 13.2 76.7
Retrieved Image 43.8 11.7 75
CHAIR(MaxLen=64)
Regular 24.4 8.9 56.6
w/o Image 25.4 7.6 61.9
Random Image 21.6 6.9 58.8
Retrieved Image 18.0 6.1 59.6
Table 8: This experiment was conducted by LLaVA-1.5. Random Image and Retrieved Image represent the contrastive images selected through random sampling and image retrieval, respectively. w/o Image represents the contrastive distribution that is constructed by removing the input image.
Figure 6: The correlation between similarity and performance. We conducted this experiment on the validation set of CHAIR. Sim-CLIP represents normalized image similarity, measured as the cosine distance between the features of two images extracted by CLIP.
C.2 Divergence Between Images
Contrastive decoding is a method that suppresses consistent features between two distributions while amplifying distinctive ones. To prevent the loss of visual information during contrastive decoding, we select an image with sufficient discrepancy to construct a contrasting distribution. Thus, while maintaining language prior consistency, the divergent visual information helps avoid erroneous debiasing caused by feature overlap. We use CLIP to retrieve images based on cosine similarity, and select the image with the lowest similarity as the contrastive image. We conduct experiments to explore the correlation between similarity and performance, as shown in Fig. 6.
We observe that the similarity between images exhibits no strong correlation with performance. Compared to regular decoding methods, images with varying similarity levels all demonstrate significant improvements. As discussed in Sec.4.1, the influence of language priors is primarily reflected in the LVLMs’ comprehension of objects, attributes, and relationships. This information exhibits limited overlap across real-world images. We hypothesize that random image selection could achieve comparable effectiveness, with experimental results presented in Tab.8. Experimental results demonstrate that random image selection can also significantly mitigate language priors in LVLMs. In practical applications, the random selection approach offers greater convenience, substantially enhancing the usability of our method.
In Tab. 8, we also attempted to construct language priors by removing visual inputs, but found this approach less effective compared to eliminating linguistic priors through cross-image consistency. We attribute this limitation to LVLMs being trained on large-scale multimodal data, where removing visual inputs disrupts the model’s intrinsic representations and leads to information distortion. Consequently, language priors in LVLMs cannot be directly constructed, which further substantiates the necessity of leveraging cross-image consistency in our method.
Figure 7: The results on the validation set of CHAIR using LLaVA-1.5. The X-axis represents the value of $\gamma$.
C.3 Threshold of Detrimental Priors
While language priors contribute to hallucinations in LVLMs, they also play a critical role in maintaining textual coherence and fluency. We employ JS divergence to distinguish between essential and detrimental priors, where essential priors exhibit high cross-image consistency, reflected in lower JS divergence scores. We establish a threshold of $\gamma=-4$ in Eq.6, based on the statistical distribution of JS divergence scores shown in Fig.5.
A comprehensive analysis is shown in Fig. 7. The experimental results align with the statistical analysis, demonstrating optimal performance at $\gamma=-4$. Another notable finding is that preserving essential priors proves meaningful, as it ensures fluent and coherent generation contexts.